OpenStudy (anonymous):

BMANN5 ur solution is here

OpenStudy (anonymous):

And yes that other guy is right i'm extremely behind on math.

OpenStudy (anonymous):

bmanny dont lose heart i dint mean to hurt u but dear , i really think that u should go to some tutor personally and should help urself, instead of finding solutions on this site

OpenStudy (anonymous):

we have to find the value of x in the following equation . \[(x/5)-12=7\] , after that we can write it as like this also \[(x/5)-(12)=7\] so we have to take Lowest common multiple of the denominatorss of x/5 and 12/1 that are 5 and 1 the LCM of 5 and 1 is 5 now divide 5 by 5 = 1 multiply 1 with x similarly divide 5 by 1= 5 multpliy 5 by 12 = 60 \[(x-60)/5 =7\] multiply both sides by 5 \[(x-60) = 7*5\] now transpose -60 to right hand side \[x= 35+60\] \[x=95\]

OpenStudy (anonymous):

i did it with another method : but the simpler method is : \[(x//5)-12 = 7\] transpose -12 to right hand side. after transposing -12 to right hand side it would become +12 ok \[x/5 = 7+12\] \[x/5 = 19\] multiply both siddes by 5 \[x = 19*5\] \[x=95\]

OpenStudy (anonymous):

see the answers are same in both the methods but it is on you which 1 u chose

OpenStudy (anonymous):

BMANNY HAVE U GOT IT NOW

OpenStudy (anonymous):

Great answer! and at least for this problem yes. Can I keep on asking questions? lol

OpenStudy (anonymous):

ofcourse u can

OpenStudy (anonymous):

yes , u can

OpenStudy (anonymous):

yes , u can

OpenStudy (anonymous):

ask it now as i am free now , after that i would not be free

OpenStudy (anonymous):

any more question ?

OpenStudy (saifoo.khan):

Kush!

OpenStudy (anonymous):

yes bhaiya